OpenStudy (anonymous):

Determine f inverse of x given the function f(x) = (x-4)^5 - 10

OpenStudy (anonymous):

I know I have to FOIL but is there a faster way to reduce the amount of work?

OpenStudy (anonymous):

You don't have to foil. Plug in x for y and then resolve for y. x = (y-4)^5 - 10 x + 10 = (y-4)^5\[y - 4 = \sqrt[5]{x + 10}\] \[y = \sqrt[5]{x + 10} + 4\]

OpenStudy (anonymous):

OK, so you took the 5th root of x and didn't foil

OpenStudy (anonymous):

@Jackienvdm

OpenStudy (anonymous):

Correct. Instead of foiling out the exponent, I took the fifth root of both sides.